# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v2

@@ III-B Analysis
$\widetilde{W}_{i}\;=\;$ $\{S_{k+1},\dots,S_{n},S_{n+i}\}\;\cup\;$
$\quad(S_{n+j}:j\neq i,j\in\mathcal{H}_{0})\;\cup(S_{n+j}:j\in\mathcal{H}_{1}).$
@@ III-C Surrogate Decision-based Attack Scheme
Step 3: Attack set selection and adversarial perturbation. Select a subset $\mathcal{A}$ from the unrejected test samples as the target, where $|\mathcal{A}|=m_{\mathcal{A}}=\lfloor\gamma(m-R)\rfloor$ with $\gamma\in(0,1]$ being the attack intensity parameter. For each $i\in\mathcal{A}$, compute
@@ Appendix A Proof of Lemma 1
$(U_{\setminus\mathcal{A}}\mid f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)\overset{d}{=}(U^{\pi}\mid f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V).$ (12)
@@
According to Assumption 1, we know that all the elements inside $U_{\setminus\mathcal{A}}$ are exchangeable given $V$. Along with the assumption that $f_{\text{attack}}$ satisfies equation 7, we have that
@@ Appendix B Proof of Lemma 2
$Q=h(U,\widetilde{U}_{\mathcal{A}},V)=((Z_{1},\dots,Z_{k}),\{\widetilde{Z}_{n+i}:i\in\mathcal{A}\cup Z_{n+i}:i\in\mathcal{H}_{1}\}).$ (14)
@@
$(U,\widetilde{U}_{\mathcal{A}},V)\overset{d}{=}(U^{\pi},\widetilde{U}_{\mathcal{A}},V).$ (15)
@@
$(U,\widetilde{U}_{\mathcal{A}},V,Q)=(U,\widetilde{U}_{\mathcal{A}},V,h(U,\widetilde{U}_{\mathcal{A}},V))\overset{d}{=}(U^{\pi},\widetilde{U}_{\mathcal{A}},V,h(U^{\pi},\widetilde{U}_{\mathcal{A}},V)).$ (16)
Since the permutation keeps those indices fixed, by the definition of $Q$, we have $h(U^{\pi},\widetilde{U}_{\mathcal{A}},V)=h(U,\widetilde{U}_{\mathcal{A}},V)=Q$. Thus
$(U,\widetilde{U}_{\mathcal{A}},V,Q)\overset{d}{=}(U^{\pi},\widetilde{U}_{\mathcal{A}},V,Q).$ (17)
Applying the score function $s$ to each $U_{i}$, we obtain
@@
Since $\pi(i)=i$ for all $i\leq k$ and $i>n+\tilde{m}_{0}$, the permutation only affects the scores
$((S_{k+1},\dots,S_{n}),S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}).$
Therefore, Eq. equation 18 implies
$(S_{k+1},\dots,S_{n}),S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q\overset{d}{=}(S_{\pi(k+1)},\dots,S_{\pi(n)}),S_{\pi(n+i)}):i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid\widetilde{U}_{\mathcal{A}},V,Q.$
Now, we justify why conditioning on $(\widetilde{U}_{\mathcal{A}},V,Q)$ is equivalent to conditioning only on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$. First, note that $Q=h(U,\widetilde{U}_{\mathcal{A}},V)$ is a deterministic function of $U$, $\widetilde{U}_{\mathcal{A}}$ and $V$. The score function $s$ depends on $\widetilde{U}_{\mathcal{A}}$ and $V$ only through this transformation $Q$. Furthermore, those conditional scores are the results of a deterministic function of $(Q,\widetilde{U}_{\mathcal{A}},V)$. Therefore, the scores $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1}\cup\mathcal{A})$ are fully determined once $\widetilde{U}_{\mathcal{A}}$, $V$ and $Q$ are fixed, and vice versa.
As a result, conditioning on $(\widetilde{U}_{\mathcal{A}},V,Q)$ is equivalent to conditioning on the non-null and attacked scores. Thus
$(\tilde{s}(Z_{k+1}),\dots,\tilde{s}(Z_{n}),\tilde{s}(Z_{n+i}):i\in\mathcal{H}_{0}\setminus\mathcal{A})$ is exchangeable conditional on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$.
@@ Appendix C Proof of Theorem 1
${\tilde{p}}_{i}=\frac{1+\sum_{j=k+1}^{n}\mathbf{1}\{\tilde{s}(Z_{j})\geq\tilde{s}(\widetilde{Z}_{n+i})\}}{n-k+1}.$
@@
Hence $p^{\prime}_{j}\leq{\tilde{p}}_{j}$ whenever ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$ is guaranteed by construction ($C_{i,j}$ is smaller or equal if ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$), and $p^{\prime}_{j}={\tilde{p}}_{j}$ if ${\tilde{p}}_{j}>{\tilde{p}}_{i}$. Condition (63) in Lemma D.6 from  [7] is satisfied for all $j\neq i$. Recall $\tilde{\tau}$ is the BH index for $({\tilde{p}}_{1},\dots,{\tilde{p}}_{m})$ and let $\tau_{i}^{\prime}:=\tau^{\prime}_{\text{BH}}$ be the BH index for $(p^{\prime}_{1},\dots,p^{\prime}_{m})$. By Lemma D.6 from  [7], we obtain
@@
$\widetilde{W}_{i}\;=\;\{S_{k+1},\dots,S_{n},S_{n+i}\}\;\cup\;(S_{n+j}:j\neq i,j\in H_{0})\;\cup\;(S_{n+j}:j\in H_{1}),$
so $\tau^{\prime}_{i}$ is $\widetilde{W}_{i}$-measurable. Hence
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\mathbb{E}\left[\frac{\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}}{\tau^{\prime}_{i}}\,\Big|\,\widetilde{W}_{i}\right]\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{1}{\tau^{\prime}_{i}}\,\mathbb{E}\left[\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}\,\Big|\,\widetilde{W}_{i}\right]\right],$
where the last equality is due to $\tau^{\prime}_{i}$ acting as a known constant inside that inner conditional. From Lemma 3, we know that $(n-k+1){\tilde{p}}_{i}$ is exactly the rank of $S_{n+i}$ among $\{S_{k+1},\dots,S_{n},S_{n+i}\}$ and ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$. As a result, $(n-k+1)\,{\tilde{p}}_{i}$ is uniform on $\{1,\dots,n-k+1\}$, independent of $\tau^{\prime}_{i}$. Thus
@@ Appendix D Direct Decision-based Attack Scheme
We start with describing our first attack scheme as follows.

### v3

@@ III-B Analysis
$\widetilde{W}_{i}\;=\;$ $(\{S_{k+1},\dots,S_{n},S_{n+i}\}\;,\;$
$\quad(S_{n+j}:j\neq i,j\in\mathcal{H}_{0})\;,(S_{n+j}:j\in\mathcal{H}_{1})\,).$
@@ III-C Surrogate Decision-based Attack Scheme
Step 3: Attack set selection and adversarial perturbation. Select a subset $\mathcal{A}$ from the unrejected test samples as the target, where $|\mathcal{A}|=m_{a}$ for some fixed number $m_{a}$. For each $i\in\mathcal{A}$, compute
@@ Appendix A Proof of Lemma 1
$(U_{\setminus\mathcal{A}}\mid f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V),V)\overset{d}{=}(U_{\setminus\mathcal{A}}^{\pi}\mid f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V),V).$ (12)
@@
According to (10), we know that all the elements inside $U_{\setminus\mathcal{A}}$ are exchangeable given $V$. Along with the assumption that $f_{\text{attack}}$ satisfies equation 7, we have that
@@ Appendix B Proof of Lemma 2
$Q=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)=((Z_{1},\dots,Z_{k}),\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{1}\cup\mathcal{H}_{0}\setminus A\}.$ (14)
@@
$(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)\overset{d}{=}(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V),$ (15)
with a slight abuse of notation where $\pi$ is defined over $1:n+\tilde{m}_{0}]$ such that $\pi(i)=i$ for $i\leq k$.
@@
$(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V,Q)=(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V,h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V))\overset{d}{=}(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V,h(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V)).$ (16)
Since the permutation keeps the first $k$ indices fixed, by the definition of $Q$, we have $h(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V)=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)=Q$. Thus
$(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V,Q)\overset{d}{=}(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V,Q).$ (17)
Applying the score function $\tilde{s}$ to each $U_{i}$ in $U_{\setminus\mathcal{A}}$, we obtain
@@
Observe that the score terms $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$ are deterministic functions of the conditional variables $\widetilde{U}_{\mathcal{A}},V,Q$. Since $\pi(i)=i$ for all $i\leq k$, the exchangeability is preserved as follows,
$(S_{k+1},\dots,S_{n},S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})\overset{d}{=}$
$(S_{\pi(k+1)},\dots,S_{\pi(n)},S_{\pi(n+i)}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A}).$
Thus $(\tilde{s}(Z_{k+1}),\dots,\tilde{s}(Z_{n}),\tilde{s}(Z_{n+i}):i\in\mathcal{H}_{0}\setminus\mathcal{A})$ is exchangeable conditional on $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$.
@@ Appendix C Proof of Theorem 1
${\tilde{p}}_{i}=\frac{1+\sum_{j=k+1}^{n}\mathbf{1}\{\tilde{s}(Z_{j})>\tilde{s}(\widetilde{Z}_{n+i})\}}{n-k+1}=\frac{1+\sum_{s\,\in\,\{S_{k+1},\dots,S_{n}\}}\mathbf{1}\{s>S_{n+i}\}}{n-k+1}.$
@@
By construction, we have $p^{\prime}_{j}\leq{\tilde{p}}_{j}$ whenever ${\tilde{p}}_{j}\leq{\tilde{p}}_{i}$, since replacing $S_{n+j}$ with a smaller score $S_{n+i}$ yields a smaller count. On the other hand, if ${\tilde{p}}_{j}>{\tilde{p}}_{i}$, it follows that $p^{\prime}_{j}={\tilde{p}}_{j}$. Also, when $i=j$, $p_{j}^{\prime}=1/(n-k+1)$ is the smallest possible value. This means that Condition (63) in Lemma D.6 from [7] is satisfied for all $(i,j)$. Recall $\tilde{\tau}$ is the BH index for $({\tilde{p}}_{1},\dots,{\tilde{p}}_{m})$ and let $\tau_{i}^{\prime}:=\tau^{\prime}_{\text{BH}}$ be the BH index for $(p^{\prime}_{1},\dots,p^{\prime}_{m})$. By Lemma D.6 from  [7], we obtain
@@
$\widetilde{W}_{i}\;=\;(\{S_{k+1},\dots,S_{n},S_{n+i}\}\;,\;(S_{n+j}:j\neq i,j\in H_{0})\;,\;(S_{n+j}:j\in H_{1})\,),$
and note that $\tau^{\prime}_{i}$ is $\widetilde{W}_{i}$-measurable. Hence
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{{\widetilde{V}}_{i}}{{{\widetilde{R}}_{m_{a}}}\lor 1}\right]$ $=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\mathbb{E}\left[\frac{\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}}{\tau^{\prime}_{i}}\,\Big|\,\widetilde{W}_{i}\right]\right]$
$=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}\setminus\mathcal{A}}\frac{1}{\tau^{\prime}_{i}}\,\mathbb{E}\left[\mathbf{1}\left\{{\tilde{p}}_{i}\leq\alpha\,(\tau^{\prime}_{i}/m)\right\}\,\Big|\,\widetilde{W}_{i}\right]\right],$
where the last equality is due to $\tau^{\prime}_{i}$ acting as a known constant inside that inner conditional. From Lemma 3, we know that $(n-k+1){\tilde{p}}_{i}$ is exactly the rank of $S_{n+i}$ among $\{S_{k+1},\dots,S_{n},S_{n+i}\}$ and ${\tilde{p}}_{i}$ is independent of $\widetilde{W}_{i}$. As a result, $(n-k+1)\,{\tilde{p}}_{i}$ is uniform over [1:nk+1][1:n-k+1], independent of $\tau^{\prime}_{i}$. Thus
@@ Appendix D Direct Decision-based Attack Scheme
We start with describing our attack scheme as follows.